Determine which of the following line, if any, are parallel or perpendicular.
Line a: 6x+2y=5 Line b: y=1/3x-4 Line c: y=-3x+5
Please explain how you arrived at your answer. Thank You.

Answer by stanbon(75887)

Determine which of the following line, if any, are parallel or perpendicular.
Line a: 6x+2y=5
y = -3x + (52)
slope = -3
-----------------------
Line b: y=1/3x-4
slope = 1/3
-------------------------

Line c: y=-3x+5
slope = -3
-------------------
lines a and b are perpendicular because the product of their
slopes is -1
-----------
lines b and c are perpendicular for the same reason
---------------
lines a and c are parallel because their slopes are equal.
===========================================================
Cheers,
Stan H.